OpenStudy (b.r):

Use the order of operations to simplify the expressions in the next two questions. 15 – 6 • 9 ÷ 2 + 10

OpenStudy (b.r):

4 • (6 + 3) + 5

jagr2713 (jagr2713):

for 1 i got -2

OpenStudy (destinycholie):

For 4x(6+3)+5 its 41

jagr2713 (jagr2713):

and 2 i got 41

OpenStudy (b.r):

How did you two get the answers exactly? I've been stuck on this forever. >.>

jagr2713 (jagr2713):

well 4*9= 36+5 = 41

OpenStudy (destinycholie):

PEMDAS Parentheses first Exponents (ie Powers and Square Roots, etc.) Multiplication and Division (left-to-right) Addition and Subtraction (left-to-right)

OpenStudy (b.r):

@jagr2713 Ah. Now I see. What about the first one?

jagr2713 (jagr2713):

15 – 6 • 9 ÷ 2 + 10 Multiply 6*9 15-54÷2+10 15-27+10 -12+10 -2

OpenStudy (b.r):

Oh. Can you help me with 2 more?

jagr2713 (jagr2713):

ok

OpenStudy (b.r):

Write the difference 30 – (–12) as a sum, then simplify. Show all steps.

jagr2713 (jagr2713):

well that's simple 30 minus negative 12 = 42

jagr2713 (jagr2713):

is that one of the answer choices

OpenStudy (b.r):

Evaluate the expression. Show all step. This is my final question. 18 + 6m for m = 2 Oh, so 42 is the answer?

jagr2713 (jagr2713):

what is 6*2?

OpenStudy (b.r):

Well for the one I've shown you above I have to write it down on my own. no answer choices. >.>

jagr2713 (jagr2713):

Yes 42 and what is 6*2

OpenStudy (b.r):

12.

jagr2713 (jagr2713):

18+12?

jagr2713 (jagr2713):

when you get that type of problem you just plug whatever equals to m in the equation

OpenStudy (b.r):

30, of course.

jagr2713 (jagr2713):

so thats your answer :D

OpenStudy (b.r):

Just like that? Oh brother. >.< I could've just done that.

jagr2713 (jagr2713):

Yea

OpenStudy (b.r):

Thanks, Btw !^-^

jagr2713 (jagr2713):

AnyTime